two cyclists start from the same point and ride on opposite directions/ One cyclist rides twice as fast as the other. On three hours they are 90 km apart. What is the rate of each cyclist?
Answer by Mathtut(3670)

distance = rate*time
lets call one cyclist d1=r1(t1)
and the other cyclist d2=r2(t2)
t1=t2=3hours
d2=90-d1
r1=2(r2)---->r1/2

so d1=r1(3)
and 90-d1=r2(3)
substitute value of d1 into 2nd equation and remember r2=r1/2
90-(3(r1))=((r1)/2)(3)---->90-3(r1)=3/2(r1)----->180-6(r1)=3(r1)--->9(r1)=180
r1=20kmph
r2=10kmph